In R^4, let U=span((1,1,0,0),(1,1,1,2)). Find u"\in"U such that ||u-(1,2,3,4)|| is as small as possible.


1
Expert's answer
2021-07-20T18:00:17-0400

"\text{We will use the normal equations for the formula of an orthogonal projection. Let}\\A=\begin{pmatrix}
 1 & 1 \\1&1\\0&1\\0&2
\end{pmatrix} and b = \begin{pmatrix}
 1 \\2\\3\\4
\end{pmatrix}
\\\text{Then u is the orthogonal projection of b onto the subspace spanned by the column }\\\text{of A and it is given by the formula:}\\u= A(A^TA)^{-1}A^Tb\\\text{We get that }\\A^Tb =\begin{pmatrix}
 1 & 1&0&0 \\1&1&1&2
\end{pmatrix}\begin{pmatrix}
 1\\2\\3\\4
\end{pmatrix}=\begin{pmatrix} 3\\14 \end{pmatrix}\\
A^TA=\begin{pmatrix}
 1 & 1&0&0 \\1&1&1&2
\end{pmatrix}\begin{pmatrix}
 1 & 1 \\1&1\\0&1\\0&2
\end{pmatrix}=\begin{pmatrix} 2&2\\2&7 \end{pmatrix}
\\(A^TA)^{-1}=\frac{1}{10}

\begin{pmatrix}
 7 & -2 \\
 -2 & 2
\end{pmatrix}\\(A^TA)^{-1}A^Tb=\frac{1}{10}

\begin{pmatrix}
 7 & -2 \\
 -2 & 2
\end{pmatrix}\begin{pmatrix}3\\14\end{pmatrix}=\frac{1}{10}

\begin{pmatrix}
 -7 \\
 22
\end{pmatrix}
\\u=A(A^TA)^{-1}A^Tb=\frac{1}{10}

\begin{pmatrix}
 1&1 \\
 1&1\\0&1\\0&2
\end{pmatrix}\begin{pmatrix}-7\\22\end{pmatrix}=\frac{1}{10}

\begin{pmatrix}
 15 \\
 15\\22\\44
\end{pmatrix}
\\\text{The answer is u=} \frac{1}{10}\begin{pmatrix}15\\15\\22\\44\end{pmatrix}"